Solve DE and Find path(time)

a) Yt+2-2yt+1+2yt=1

y0=1 y1=4

b) Yt+2-yt+1+1/4yt=1

Y0=4 , Y1=7
Answer #1

a) Solution ; y(t+2) - 2 y(t+1) + 2 y(«):1, . 7(0) = 1, y(i)=4. The equation can be written as, (DP-2D +2 ) y = 1. where , heb) y(++2) - y(t+1) +57(+)=1 , 460)=4, 4(1)-7. The equation can be written as, 1 (0²-D + 4) y =1., where af al =D Now, let y-e